# DRDT3: Diffusion-Refined Decision Test-Time Training Model

### v1

@@ III-A2 Coarse Action Prediction
5:     Compute $\mathcal{L}_{dt3}$ by Equation 16
@@
8:     Compute $\mathcal{L}_{diff}$ by Equation 15
9:      Update policy by minimizing $\mathcal{L}_{DRDT3}=\mathcal{L}_{diff}(\theta)+\zeta\mathcal{L}_{dt3}(\theta)$
@@
III-B Conditional Diffusion Policy
We adopt DDPM [18], a well-known generative model, as the enhancement module to refine the coarse actions predicted from DT3. To achieve this, we extend DDPM to a conditional diffusion model and employ coarse action $\widetilde{\bm{a}}$ predicted from DT3 as a prior, as depicted in Figure 2. Consequently, the final action is sampled through a reversed process:
$\begin{array}[]{c}\pi_{\theta}(\bm{a}|\widetilde{\bm{a}}):=p_{\theta}\left(\bm{a}^{0:N}|\widetilde{\bm{a}}\right)\\ :=p\left(\bm{a}^{N}\right)\prod_{i=1}^{N}p_{\theta}\left(\bm{a}^{i-1}\mid\bm{a}^{i},\widetilde{\bm{a}}\right),\end{array}$ (11)
where the final denoised sample $\bm{a}^{0}$ denotes the sampled action to be implemented. It is worth noting that we use subscripts $t\in\{1,...,T\}$ to denote timesteps in trajectories while using superscripts $i\in\{1,...,N\}$ to represent diffusion timesteps. We formulate $p_{\theta}\left(\bm{a}^{i-1}\mid\bm{a}^{i},\widetilde{\bm{a}}\right)$ using a Gaussian transition
where the covariance matrix is set as $\mathbf{\Sigma}_{\theta}\left(\bm{a}^{i},\widetilde{\bm{a}},i\right)=\beta^{i}\mathbf{I}$. Following DDPM [18], we can infer the mean as
@@ III-B Conditional Diffusion Policy
2:for $i=N,...1$ do
3:      Predict coarse action representation $\widetilde{\bm{a}}$ with the DT3 module
@@
5:     $\bm{a}^{i-1}=\frac{1}{\sqrt{\alpha^{i}}}\left(\bm{a}^{i}-\frac{1-\alpha^{i}}{\sqrt{1-\bar{\alpha}^{i}}}\bm{\epsilon}_{\theta}\left(\bm{a}^{i},\widetilde{\bm{a}},i\right)\right)+\beta^{i}\bm{\epsilon}$
@@
7:Return $\bm{a}^{0}$ as the selected action
@@
where $\bm{\epsilon}\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)$ when $i>1$ while $\bm{\epsilon}=0$ when $i=1$ for better sample quality [18]. We choose to predict the noise $\bm{\epsilon}$ using neural networks. Therefore, according to DDPM [18], the noise approximator $\bm{\epsilon}_{\theta}$ can be optimized with a simplified loss function
@@
III-C Gated MLP Noise Approximator
To fully integrate coarse action predictions into the diffusion model and improve noise prediction accuracy, we introduce a gated MLP noise approximator, as depicted in the right part of Figure 3. The gated MLP noise approximator utilizes adaptive layer normalization (adaLN) [27] to capture condition-specific information. It specifically learns the scale and shift parameters, $\gamma$ and $\beta$, for layer normalization, as well as a scaling parameter $\alpha$ based on the concatenated condition of the diffusion timestep and coarse action prediction. Additionally, we employ a gated MLP block [28, 29] to extract multi-granularity information from the input, enhancing the model’s noise reconstruction capabilities. This architecture splits the output from the Scale & Shift layer into two separate pathways, each expanding the feature dimensionality by a factor of $M$ using a linear layer. The GeLU [30] non-linearity is applied in one branch, after which the two streams are combined through element-wise multiplication. A final linear layer reduces back the dimensionality, and a residual connection is incorporated to facilitate better gradient flow.
III-D Unified optimization objective
The diffusion loss $\mathcal{L}_{diff}(\theta)$ facilitates refinement for the coarse action representation $\widetilde{\bm{a}}$ and behavior cloning towards the dataset, while it fails to constrain the DT3 model, potentially leading to deviations in coarse action predictions from the data distribution. To alleviate such issues and jointly optimize DT3 and the diffusion model, we propose a unified optimization objective. This involves introducing a DT3 loss as additional guidance during the training of the diffusion model
@@ III-D Unified optimization objective
where $\bm{a}_{-K}$ is the ground truth of the latest $K$-step actions, and $\bm{a}_{\text{max}}$, the maximum action in the action space, serves as a scaling factor to render the $L_{1}$ loss dimensionless. It is worth noting that the diffusion model only conditions on the last term $\widetilde{\bm{a}}$ within the sequence $\widetilde{\bm{a}}_{-K,t}$, whereas the DT3 loss is computed for the entire sequence. The final unified loss for DRDT3 is a linear combination of the DT3 loss (action representation term) and diffusion loss (action refinement term)
@@
where $\zeta$ is a loss coefficient for balancing the two loss terms. Algorithm 1 and 2 illustrate the training and inference procedures of DRDT3.
TABLE I: Hyperparameter settings for denoiser network.
Hyperparameter Value
learning rate 0.0003
optimizer Adam
batch size 2048
number of gradient update steps per epoch 2000
number of training epochs 50
number of Attention TTT blocks 1
TTT layer type linear
number of hidden layers in TTT layer 6
embedding dimension 128
diffusion timestep embedding dimension 16
gated MLP width 384
gated MLP activation GeLU
number of diffusion steps 5
$\beta_{\min}$ 0.1
$\beta_{\max}$ 10.0
context length $K$ 6
RTG scale factor 1.7
loss coefficient $\zeta$ 0.2
Figure 4: Gym-MuJoCo locomotion and AntMaze tasks on D4RL benchmark.
III-E Implementation Details
When implementing our proposed DRDT3, we train it for 50 epochs with 2000 gradient updates per epoch. The learning rate and batch size are designated as 0.0003 and 2048, respectively. To proceed with historical subtrajectories with DT3, we set the context length as 6. The Attention TTT block used in the DT3 module consists of 1-layer self-attention and 1-layer TTT with embedding dimensions of 128. We use a linear layer to output a deterministic coarse representation of the action. To enhance computational efficiency without sacrificing performance, the number of diffusion timesteps is set to 5. We set the variance schedule according to Variance Preserving (VP) SDE [31]:
@@ III-E Implementation Details
Most of the hyperparameters are selected through the Optuna hyperparameter optimization framework [32].
IV Experiments
We conduct experiments to evaluate our proposed DRDT3 on the commonly-used D4RL benchmark [23] using an AMD Ryzen 7 7700X 8-Core Processor with a single NVIDIA GeForce RTX 4080 GPU.
IV-A Experimental Settings
IV-A1 Tasks and Datasets
As shown in Figure 4, we adopt four Mujoco locomotion tasks from Gym, which are HalfCheetah, Hopper, Walker2D and Ant, and a goal-reaching task, AntMaze. For the Gym tasks, dense rewards are provided, and we use three different dataset settings from D4RL: Medium (m), Medium-Replay (mr), and Medium-Expert (me). Specifically, Medium datasets contain one million samples collected using a behavior policy which achieves $1/3$ scores of an expert policy; Medium-Reply datasets contain the replay buffer during training a policy till reaching the Medium score; Medium-Expert datasets consist of two million samples evenly generated from a medium agent and an expert agent. The AntMaze task aims to move an ant robot to reach a target location with sparse rewards (1 for reaching the goal otherwise 0). Following experimental settings in a DT variant [33], we adopt umaze and umaze-diverse datasets for evaluation.
IV-A2 Baselines
@@
  • TD3+BC [34]: An offline RL algorithm that adapts the TD3 method [35] by incorporating a behavior cloning regularization term into the policy objective. This modification aims to prevent the policy from generating actions that deviate significantly from the offline dataset distribution.
@@
  • Conservative Q-Learning (CQL) [36]: An offline RL algorithm that penalizes the Q-values associated with out-of-distribution actions, thereby encouraging more conservative Q-function estimates.
@@
  • Implicit Q-Learning (IQL) [12]: IQL approximates the upper expectile of Q-value distributions, allowing it to learn policies focusing on actions within the dataset distribution.
  • Diffuser [21]: A diffusion-based algorithm that adopts diffusion models as a planner for generating trajectories in the form of model-based offline RL.
  • DT [14]: DT formulates the policy learning as a sequence modeling problem by autoregressively generating trajectories using GPT models.
  • Contrastive DT (ConDT) [37]: A DT variant that harnesses the power of contrastive representation learning and return-dependent transformations to cluster the input embeddings based on their associated returns.
  • Q-learning DT (QDT) [10]: A DT variant that improves the stitching ability of DT by introducing dynamic programing-derived value function.
  • Elastic DT (EDT) [17]: A DT variant that dynamically adjusts the length of the input context according to the quality of the previous trajectory.
The best scores reported in papers of these baselines are adopted. For ConDT, we calculate the normalized score according to the returns reported in [37].
@@
TABLE II: Normalized scores of different offline RL methods on Gym and AntMaze tasks from D4RL benchmark. Results of DRDT3 represent the mean and variance across three seeds. The best scores are highlighted in bold.
Dataset BC TD3BC CQL IQL Diffuser DT ConDT QDT EDT DRDT3 (ours)
halfcheetah-m 42.6 48.3 44.0 47.4 42.8 42.6 43.0 42.3 42.5 44.1±1.3
hopper-m 52.9 59.3 58.5 66.3 74.3 67.6 74.5 66.5 63.5 92.7±8.8
walker2d-m 75.3 83.7 72.5 78.3 79.6 74 72.8 67.1 72.8 84.2±1.8
halfcheetah-mr 36.6 44.6 45.5 44.2 37.7 36.6 40.9 35.6 37.8 42.3±0.5
hopper-mr 18.1 60.9 95.0 94.7 93.6 82.7 95.1 52.1 89.0 92.7±2.1
walker2d-mr 26.0 81.8 77.2 73.9 70.6 66.6 72.5 58.2 74.8 85.0±1.8
halfcheetah-me 55.2 90.7 91.6 86.7 88.9 86.8 92.6 79.0 89.1 94.2±2.1
hopper-me 52.5 98.0 105.4 91.5 103.3 107.6 110.3 94.2 108.7 112.5±1.5
walker2d-me 107.5 110.1 108.8 109.6 106.9 108.1 109.1 101.7 106.2 104.5±6.4
Average 51.9 75.3 77.6 77 77.5 74.7 79.0 66.3 76.0 83.6
ant-m - - - 99.9 - 93.6 - - 97.9 103.3±5.8
ant-mr - - - 91.2 - 89.1 - - 92.0 96.7±5.1
Average - - - 95.6 - 91.4 - - 95.0 100.0
antmaze-umaze 54.6 78.6 74.0 87.5 - 59.2 - 67.2 - 76.6±1.9
antmaze-umaze-diverse 45.6 71.4 84.0 62.2 - 53 - 62.1 - 82.5±6.3
Average 50.1 75.0 79.0 74.9 - 56.1 - 64.7 - 79.6
IV-B Performance Comparison
@@
Table II presents the normalized scores [23] of our DRDT3 and baselines on several tasks from the D4RL benchmark. As we can see, DRDT3 significantly improves the average scores on both Gym and AntMaze tasks. Specifically, DRDT3 outperforms other offline RL methods including the state-of-the-art DT variants on 7 out of 11 Gym tasks. For the other 4 Gym tasks, DRDT3 is still comparable over the best baselines. Particularly, the Medium-Expert dataset is not reported for the Ant task given the lack of comparing baselines evaluated on it. For AntMaze tasks, DRDT3 shows comparable performance over the best baselines for each single task while achieving the best overall results. When compared with the vanilla DT, our DRDT3 presents significant improvements on 12 out of 13 tasks, especially on tasks using Medium and Medium-Replay datasets with sub-optimal trajectories, which demonstrates the enhanced stitching ability of our DRDT3 over DT.
IV-C Ablation Study
We implement ablation studies for some components to demonstrate their effectiveness.
IV-C1 Ablation on Sequence Model
We compare our DRDT3 framework with two variants: one where the sequence model is replaced by a GPT-2 model [16] and another where the TTT layer is replaced by a MAMBA layer [38]. The GPT-2 model relies entirely on the attention mechanism, while MAMBA is a selective state space model based on the RNN structure, where parameters are dynamically dependent on the inputs. For the two variants w/ GPT-2 and w/ Attention MAMBA, we use one block per model as our DRDT3. To ensure a fair comparison, we adjust the number of layers in both the GPT-2 and MAMBA variants to match the model size of our DRDT3. Specifically, the total number of learnable parameters for our DRDT3, the variant w/ GPT-2, and the variant w/ Attention MAMBA are 235K, 218K, and 329K, respectively. We present the structures of three noise approximations adopted for comparison in Figure 6.
As shown in Figure 5, our DRDT3 with Attention TTT blocks outperforms the other variants across three Medium-Expert tasks. The variant with Attention MAMBA shows slight performance degradation and lacks the stability of DRDT3. The variant with GPT-2 performs the worst across all tasks, which might be attributed to typical reliance on larger model sizes of GPT-2. When constrained to a similar size as the other two variants, GPT-2 experiences performance degradation. These results indicate that our proposed DT3 module with Attention TTT blocks enhances performance without increasing model size. It is worth noting that our DRDT3 does not show improved computation efficiency over other variants, but the final performance is promising. This may be due to the use of short-range context as the sequence input, whereas the MAMBA and TTT layers are optimized for more efficient long-sequence processing.
IV-C2 Ablation on Diffusion
To test whether the diffusion module plays a significant role in action refinement and verify the effectiveness of our DT3 module, we train a DT3 model alone with the same settings as DRDT3 by minimizing the $\mathcal{L}_{dt3}$ in Equation 16. Table III and IV show the performance comparison between DT and our DT3 model on Gym and AntMaze tasks. As we can see, our proposed DT3 trained independently can outperform DT on 6 out of 9 Gym locomotion tasks and all on all AntMaze tasks. This demonstrates the superiority of our proposed DT3 with the Attention TTT blocks over the vanilla DT with only the attention mechanism. Furthermore, our proposed DRDT3, which enhances DT3 through a diffusion module, further boosts performance across all tasks with different datasets, achieving average improvements of 10.3%, 7.6%, and 38.0% on Gym locomotion, Ant, and AntMaze tasks, respectively. This demonstrates the substantial impact of incorporating diffusion refinement in DRDT3.
@@ IV-C2 Ablation on Diffusion
TABLE III: Performance comparison between DT, DT3 and DRDT3 on Gym locomotion tasks. The underlined values indicate performance better than DT and bold values are the best results. ’hc’: halfcheetah; ’hp’: hopper; ’wk’: walker2d.
@@
TABLE IV: Performance comparison between DT3 and DT on Ant and AntMaze tasks.

### v2

@@ 4.1.2 Coarse Action Prediction
4.2 Conditional Diffusion Policy
We adopt DDPM (Ho et al., 2020), a well-known generative model, as an enhancement module to refine the coarse actions predicted from DT3. To achieve this, we extend DDPM to a conditional diffusion model and employ coarse action $\widetilde{\bm{a}}$ predicted from DT3 as a prior, as depicted in Figure 1. Consequently, the final action is sampled through a reversed process:
[…]
where the covariance matrix is set as $\mathbf{\Sigma}_{\theta}\left(\bm{a}^{i},\widetilde{\bm{a}},i\right)=\beta^{i}\mathbf{I}$. Following DDPM (Ho et al., 2020), we can infer the mean as
@@ 4.2 Conditional Diffusion Policy
5:  Compute $\mathcal{L}_{dt3}$ by Equation 16 $\triangleright$ DT loss
@@
8:  Compute $\mathcal{L}_{diff}$ by Equation 15 $\triangleright$ Diffusion loss
9:  Update policy by minimizing $\mathcal{L}_{DRDT3}=\mathcal{L}_{diff}(\theta)+\zeta\mathcal{L}_{dt3}(\theta)$ $\triangleright$ Unified objective
@@
2:for i=N,1i=N,...1 do
3:  Predict coarse action representation $\widetilde{\bm{a}}$ with the DT3 module $\triangleright$ Coarse action prediction
@@
5:  $\bm{a}^{i-1}=\frac{1}{\sqrt{\alpha^{i}}}\left(\bm{a}^{i}-\frac{1-\alpha^{i}}{\sqrt{1-\bar{\alpha}^{i}}}\bm{\epsilon}_{\theta}\left(\bm{a}^{i},\widetilde{\bm{a}},i\right)\right)+\beta^{i}\bm{\epsilon}$ $\triangleright$ Action refinement
@@
7:Return $\bm{a}^{0}$ as the selected action $\triangleright$ Final action
@@
where $\bm{\epsilon}\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)$ when $i>1$ while $\bm{\epsilon}=0$ when $i=1$ for better sample quality (Ho et al., 2020). We choose to predict the noise $\bm{\epsilon}$ using neural networks. Therefore, according to DDPM (Ho et al., 2020), the noise approximator $\bm{\epsilon}_{\theta}$ can be optimized with a simplified loss function
@@
4.3 Gated MLP Noise Approximator
To fully integrate coarse action predictions into the diffusion model and improve noise prediction accuracy, we introduce a gated MLP noise approximator, as depicted in the right part of Figure 2. The gated MLP noise approximator utilizes adaptive layer normalization (adaLN) (Peebles & Xie, 2023) to capture condition-specific information. It specifically learns the scale and shift parameters, $\gamma$ and $\beta$, for layer normalization, as well as a scaling parameter $\alpha$ based on the concatenated condition of the diffusion timestep and coarse action prediction. Additionally, we employ a gated MLP block (Dauphin et al., 2017; De et al., 2024) to extract multi-granularity information from the input, enhancing the model’s noise reconstruction capabilities. This architecture splits the output from the Scale & Shift layer into two separate pathways, each expanding the feature dimensionality by a factor of $M$ using a linear layer. The GeLU (Hendrycks & Gimpel, 2016) non-linearity is applied in one branch, after which the two streams are combined through element-wise multiplication. A final linear layer reduces back the dimensionality, and a residual connection is incorporated to facilitate better gradient flow.
4.4 Unified optimization objective
The diffusion loss $\mathcal{L}_{diff}(\theta)$ facilitates refinement for the coarse action representation $\widetilde{\bm{a}}$ and behavior cloning towards the dataset, while it fails to constrain the DT3 module, potentially leading to deviations in coarse action predictions from the data distribution. To alleviate such issues and jointly optimize DT3 and the diffusion model, we propose a unified optimization objective. This involves introducing a DT3 loss as additional guidance during the training of the diffusion model
@@ 4.4 Unified optimization objective
where $\bm{a}_{-K}$ is the ground truth of the latest $K$-step actions, and $\bm{a}_{\text{max}}$, the maximum action in the action space, serves as a scaling factor to render the $L_{1}$ loss dimensionless. We choose $L_{1}$ loss over $L_{2}$ loss since the former is more robust to outliers and provides stable gradients within the unified objective in Equation 17, without weakening the influence of the diffusion loss. The directional information provided by $L_{1}$ loss further enhances the effectiveness of coarse-to-fine refinement. It is worth noting that the diffusion model only conditions on the last term $\widetilde{\bm{a}}$ within the sequence $\widetilde{\bm{a}}_{-K,t}$, whereas the DT3 loss is computed for the entire sequence.
The final unified loss for DRDT3 is a linear combination of the DT3 loss (action representation term) and diffusion loss (action refinement term)
@@
where $\zeta$ is a loss coefficient for balancing the two loss terms. Note that the DT3 and diffusion modules are jointly trained in a single stage by minimizing $\mathcal{L}_{\text{DRDT3}}$.
4.5 Implementation Details
Algorithm 1 and 2 illustrate the training and inference procedures of DRDT3. When implementing our proposed DRDT3, we train it for 50 epochs with 2000 gradient updates per epoch. The learning rate and batch size are designated as 0.0003 and 2048, respectively. To proceed with historical subtrajectories with DT3 module, we set the context length as 6. The Attention TTT block used in the DT3 module consists of 1-layer self-attention and 1-layer TTT with embedding dimensions of 128. We use a linear layer to output a deterministic coarse representation of the action. To enhance computational efficiency without sacrificing performance, the number of diffusion timesteps is set to 5. We set the variance schedule according to Variance Preserving (VP) SDE (Xiao et al., 2021):
@@ 4.5 Implementation Details
Most of the hyperparameters are selected through the Optuna hyperparameter optimization framework (Akiba et al., 2019).
5 Experiments
We conduct experiments to evaluate our proposed DRDT3 on the commonly used D4RL benchmark (Fu et al., 2020) using an AMD Ryzen 7 7700X 8-Core Processor with a single NVIDIA GeForce RTX 4080 GPU.
5.1 Experimental Settings
5.1.1 Tasks and Datasets
We adopt four Mujoco locomotion tasks from Gym, which are HalfCheetah, Hopper, Walker2D, and Ant, and a goal-reaching task, AntMaze. We further evaluate our methods on more challenging tasks, including the Pen and Door manipulation tasks from the Adroit benchmark, which involve a 24-DoF robot hand, as well as Humanoid locomotion tasks from the Gym environment featuring 348-dimensional observations and 17-dimensional actions. For the Gym tasks, dense rewards are provided, and we use three different dataset settings from D4RL (Fu et al., 2020): Medium (m), Medium-Replay (mr), and Medium-Expert (me). Specifically, Medium datasets contain one million samples collected using a behavior policy that achieves $1/3$ scores of an expert policy; Medium-Reply datasets contain the replay buffer during training a policy till reaching the Medium score; Medium-Expert datasets consist of two million samples evenly generated from a medium agent and an expert agent. The AntMaze task aims to move an ant robot to reach a target location with sparse rewards (1 for reaching the goal, otherwise 0). Following experimental settings in a DT variant (Zheng et al., 2022), we adopt the umaze and umaze-diverse datasets for evaluation. For the Pen and Door tasks, we use the human and cloned datasets from D4RL. For the Humanoid tasks, we adopt three offline datasets provided by Bhargava et al. (2023): Medium, Medium-Expert, and Expert.
5.1.2 Baselines
@@
TD3+BC (Fujimoto & Gu, 2021): It adapts the TD3 method (Fujimoto et al., 2018) by incorporating a behavior cloning regularization term into the policy objective. This modification aims to prevent the policy from generating actions that deviate significantly from the offline dataset distribution.
@@
Conservative Q-Learning (CQL) (Kumar et al., 2020): It penalizes the Q-values associated with out-of-distribution actions, thereby encouraging more conservative Q-function estimates.
@@
Implicit Q-Learning (IQL) (Kostrikov et al., 2021): IQL approximates the upper expectile of Q-value distributions, allowing it to learn policies focusing on actions within the dataset distribution.
@@
Table 1: Normalized scores of different offline RL methods on Gym and AntMaze tasks from D4RL benchmark. Results of DRDT3 represent the mean and variance across three seeds. The best scores are highlighted in bold.
Dataset Environment BC TD3+BC CQL IQL DT ConDT QDT EDT DRDT3 (ours)
Medium HalfCheetah 42.6 48.3 44.0 47.4 42.6 43.0 42.3 42.5 44.1±1.3
Hopper 52.9 59.3 58.5 66.3 67.6 74.5 66.5 63.5 92.7±8.8
Walker2d 75.3 83.7 72.5 78.3 74.0 72.8 67.1 72.8 84.2±1.8
Medium-Replay HalfCheetah 36.6 44.6 45.5 44.2 36.6 40.9 35.6 37.8 42.3±0.5
Hopper 18.1 60.9 95.0 94.7 82.7 95.1 52.1 89.0 92.7±2.1
Walker2d 26.0 81.8 77.2 73.9 66.6 72.5 58.2 74.8 85.0±1.8
Medium-Expert HalfCheetah 55.2 90.7 91.6 86.7 86.8 92.6 79.0 89.1 94.2±2.1
Hopper 52.5 98.0 105.4 91.5 107.6 110.3 94.2 108.7 112.5±1.5
Walker2d 107.5 110.1 108.8 109.6 108.1 109.1 101.7 106.2 104.5±6.4
Average 51.9 75.3 77.6 77.0 74.7 79.0 66.3 76.0 83.6
Medium Ant - - - 99.9 93.6 - - 97.9 103.3±5.8
Medium-Replay Ant - - - 91.2 89.1 - - 92.0 96.7±5.1
Average - - - 95.6 91.4 - - 95.0 100.0
Umaze Antmaze 54.6 78.6 74.0 87.5 59.2 - 67.2 - 76.6±1.9
Umaze-Diverse Antmaze 45.6 71.4 84.0 62.2 53.0 - 62.1 - 82.5±6.3
Average 50.1 75.0 79.0 74.9 56.1 - 64.7 - 79.6
DT-based algorithms:
DT (Chen et al., 2021): DT formulates the policy learning as a sequence modelling problem by autoregressively generating trajectories using GPT models.
Contrastive DT (ConDT) (Konan et al., 2023): It harnesses the power of contrastive representation learning and return-dependent transformations to cluster the input embeddings based on their associated returns.
Q-learning DT (QDT) (Yamagata et al., 2023): It improves the stitching ability of DT by introducing dynamic programming-derived value function.
Elastic DT (EDT) (Wu et al., 2023): It dynamically adjusts the length of the input context according to the quality of the previous trajectory.
Diffusion model-based RL methods:
Synther (Lu et al., 2023): A diffusion model-based data augmentation method that directly samples synthetic offline data. We use TD3+BC as the evaluation algorithm on the augmented datasets.
Diffuser (Janner et al., 2022): A diffusion-based algorithm that adopts diffusion models as a planner for generating trajectories in the form of model-based offline RL.
Decision Diffuser (DD) (Ajay et al., 2023): A sequence modeling method that samples future states with a diffusion model and infers actions via inverse dynamics.
AdaptDiffuser (Liang et al., 2023): It improves the diffusion model-based planner by self-evolving with filtered high-quality synthetic demonstrations.
EDP (Kang et al., 2023): A diffusion-based policy gradient method that accelerates training by approximating actions from corrupted samples in a single step.
The best scores reported in papers of these baselines are adopted. For ConDT, we calculate the normalized score according to the returns reported in (Konan et al., 2023). For diffusion-based methods, we adopt reproduced results in CleanDiffuser (Dong et al., 2024).
Table 2: Performance comparison of DRDT3 and diffusion model-based methods on Gym tasks. While DRDT3 does not surpass the baselines on individual tasks, it achieves the highest average normalized score across all tasks.
Dataset Environment SynthER Diffuser $D$ AdaptDiffuser EDP DRDT3 (ours)
[…]
Average 69.4 78.2 83.2 82.4 82.4 83.6
Table 3: Performance comparison on challenging robot manipulation tasks, i.e., Pen and Door, and the Humanoid task with high-dimensional observations.
Tasks BC CQL DT DRDT3 (ours)
[…]
Average 16.1 18.2 31.4 42.9
[…]
Average 17.0 55.8 43.2 55.7
5.2 Performance Comparison
@@
Table 1 presents the normalized scores (Fu et al., 2020) of our DRDT3, classic offline RL and DT-based baselines on several tasks from the D4RL benchmark. As we can see, DRDT3 significantly improves the average scores on both Gym and AntMaze tasks. Specifically, DRDT3 outperforms other offline RL methods, including the state-of-the-art DT variants, on 7 out of 11 Gym tasks. For the other 4 Gym tasks, DRDT3 is still comparable over the best baselines. Particularly, the Medium-Expert dataset is not reported for the Ant task, given the lack of comparing baselines evaluated on it. For AntMaze tasks, DRDT3 shows comparable performance over the best baselines for each single task while achieving the best overall results. When compared with the vanilla DT, our DRDT3 presents significant improvements on 12 out of 13 tasks, especially on tasks using Medium and Medium-Replay datasets with sub-optimal trajectories, which demonstrates the enhanced stitching ability of our DRDT3 over DT.
@@
We further evaluate DRDT3 by comparing it with recent diffusion model-based methods. As shown in Table 2, DRDT3 achieves superior performance on only one individual task, yet maintains consistently strong results across all tasks, with only a small gap to the best-performing method. Overall, it attains the highest average normalized score among all diffusion-based methods, underscoring its robustness. This outcome may stem from the fact that DRDT3 is inherently constrained by the performance of its underlying DT module, as it directly refines DT-generated actions without incorporating dynamic programming. Consequently, DRDT3’s performance is steady rather than outstanding on individual tasks.
Table 3 presents the evaluation performance of DRDT3 on more challenging tasks, including the Pen and Door tasks from the Adroit benchmark and the Humanoid task from Gym. We report results for BC, CQL, and DT as baselines, since other methods are rarely evaluated on these tasks. As shown, DRDT3 consistently outperform the baseline methods on most Pen and Door datasets, with DRDT3 achieving an average improvement of 36.6% over DT. For the Humanoid task, DRDT3 attains performance comparable to the best baseline, CQL, while still delivering a 28.9% improvement over DT. The underperformance observed on certain tasks likely stems from an inherent limitation of sequence modeling methods, which lack dynamic programming capabilities to effectively extrapolate toward higher performance.
Our primary goal, however, is to demonstrate that diffusion models can effectively refine DT-generated actions and enhance DT’s trajectory-stitching capability—an improvement our experiments confirm across a wide range of DT-based baselines and evaluation tasks. In particular, DRDT3 achieves average improvements of 11.9%, 9.4%, 41.9%, 36.6%, and 28.9% over DT on Gym locomotion, Ant, AntMaze, Adroit, and Humanoid tasks, respectively.
Table 4: Performance comparison between DT, DT3 and DRDT3 on Gym locomotion tasks. The underlined values indicate performance better than DT, and bold values are the best results. ‘hc’: halfcheetah; ‘hp’: hopper; ‘wk’: walker2d.
@@
Table 5: Performance comparison between DT3 and DT on Ant and AntMaze tasks.